# Everything everywhere all at once, a probability-based enhanced sampling approach to rare events

### v1

@@ IV Conclusions
All the authors equally contributed to the manuscript by conceptualizing the project, developing the theoretical methodologies, and participating in the writing of the manuscript.
Acknowledgements.
The Authors are grateful to Luigi Bonati, Andrea Rizzi, Jintu Zhang, Umberto Raucci, and Francesco Mambretti for the useful discussions and feedback about this manuscript.
@@ Acknowledgements.
To improve the numerical stability and efficacy of the Kolmogorov bias, it is convenient to reformulate its calculation (Eq. 6) in terms of the committor-related $z$ collective variable as it allows lifting the numerical problems that may arise from the direct use of $q$ (see Fig. S1). Here, we show that this is formally equivalent to expressing as a functional of $q$. In the main text, we shown that $q$ is obtained applying a sigmoid-like switching function $\sigma$ to $z$
@@
Figure S1: Visualization of the numerical issue related to the committor function on a toy double-well potential for which an analytical solution can be obtained.
@@
To have a direct comparison with the reference numerical result ${K}_{m}=4.18$, which is taken from Ref. 1, the reported ${K}_{m}$ values are computed on the same ideal dataset described in Ref. 1, which was obtained from a homogeneous grid (that is, 200*200 evenly distributed points) in the relevant part of the Cartesian space (that is, -1.4<x<1.1 and -0.25<y<2.0 ) assigned with their analytical Boltzmann weights.
Table S1: Summary of the iterative procedure for Müller-Brown potential.
@@
Table S2: Summary of the iterative procedure for Alanine.
@@
Figure S3: Sampling at different iterations on the procedure on Alanine. Top row: Scatter plot of the sampled points in the $\varphi\psi$ space. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red.. Middle row: Time series of the $\varphi$ value from a trajectory started from state A. Bottom row: Time series of the $\varphi$ value from a trajectory started from state B.
@@
Table S3: Summary of the iterative procedure for double path potential.
@@
Figure S4: Sampling at different iterations of the procedure on the double path potential. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red. The 0.5 value of the committor estimate at that iteration is provided as a black line.
@@
Table S4: Summary of the iterative procedure for chignolin.
@@
Comparison between different descriptor sets In Ref. 1, we trained the committor model using as inputs the 45 distances between $\alpha$-C atoms, whereas here, we used the same set of 210 distances identified in Ref. 19. To compare the two descriptor sets, we optimized the committor model on the same dataset using the two sets, obtaining a $K_{m}$ value of 2.17 for the 210 distances and 18.5 for the 45 distances. This result clearly shows how the greater variational flexibility introduced by side-chain information allows for capturing more features of the folding process.
@@
Table S5: Summary of LASSO analysis for chignolin.
@@
Figure S5: Free energy estimates for chignolin projected along $z$ and $q$ obtained averaging on 6 independent simulations: 3 started from the folded state and 3 from the unfolded one. The average value and the relative standard deviation are reported as a solid line and a shaded region, respectively.
@@
The funnel restraint In our simulations, we used a funnel restraint Limongelli, Bonomi, and Parrinello (2013) equivalent to the one previously employed by Refs.37; 67; 68 on the same system. Here, we summarize the details of such a restraint, while more details can be found in the original works. The funnel limits the space available to the ligand in the unbound state U by confining it to a cylindrical volume above the binding site. As the ligand approaches the binding site, the funnel restraint becomes wider so that its presence does not affect the binding process itself. Having aligned with PLUMED the system to a reference configuration where the binding axis is found along the vertical axis, we define $h$ as the projection on the binding axis of the center of the carbon atoms of each ligand and $r$ as its radial component. When $h>10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel surface is a cylinder with radius $R_{cyl}=2$\mathrm{\SIUnitSymbolAngstrom}$$ with its axis along the vertical direction. When $h<10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel opens into an umbrella-like shape with a 45 degree angle whose surface is defined by $r=12-h$. The force that, for a displacement x, pushes the ligand away from the funnel’s surface is harmonic $-k_{F}x$ with $kF=20$ kJ/mol Å$^{-2}$. A further harmonic restraint is applied on $h$ to prevent the ligand from getting too far from the host, reaching the upper boundary of the simulation box. The corresponding force is $-k_{U}(h-18)$ for $h>18$\mathrm{\SIUnitSymbolAngstrom}$$ and $k_{U}=40$ kJ/mol Å$^{-2}$.
@@
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref. 37, 8 with respect to 2.5Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)), and a set of distances between the same 4 ligand atoms and the binding pocket, instead of the $h$ projection only, as inputs of a neural network (NN) with architecture [16, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for $\sim$30000 epochs. The $\alpha$ hyperparameter in the loss function was set to 1. In the loss function, we assigned to the virtual points along the binding axis used for the computation of water coordination numbers an infinite mass, i.e., $10^{6}$. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S6 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S6: Summary of the iterative procedure for calixarene.
@@
Figure S6: Descriptors ranking for the committor model of calixarene trained using the descriptors listed in the computational details.
@@
Figure S7: Free energy estimates for calixarene projected along the projection of the ligand molecule on the binding axis $h$, $z$ and $q$.
Figure S8: Two-dimensional free energy surface for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
@@
Figure S9: Two-dimensional Kolmogorov Ensemble projection for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).

### v4

@@ Acknowledgements.
All the authors equally contributed to the manuscript by conceptualizing the project, developing the theoretical methodologies, and participating in the writing of the manuscript. Specifically, E.T. and P.K. developed the code for the training of the committor function and conducted the computational simulations.
Competing interests statement
The authors declare no competing interests.
@@
To improve the numerical stability and efficacy of the Kolmogorov bias, it is convenient to reformulate its calculation (Eq. 6) in terms of the committor-related $z$ collective variable as it allows lifting the numerical problems that may arise from the direct use of $q$ (see Figure S1). Here, we show that this is formally equivalent to expressing as a functional of $q$. In the main text, we shown that $q$ is obtained applying a sigmoid-like switching function $\sigma$ to $z$
@@
Figure S1: Numerical issues of the committor function Visualization of the numerical issue related to the committor function on a toy double-well potential along one dimension x for which an analytical solution can be obtained.
@@
To have a direct comparison with the reference numerical result ${K}_{m}=4.18$, which is taken from Ref. 3, the reported ${K}_{m}$ values are computed on the same ideal dataset described in Ref. 3, which was obtained from a homogeneous grid (that is, 200*200 evenly distributed points) in the relevant part of the Cartesian space (that is, -1.4<x<1.1 and -0.25<y<2.0 ) assigned with their analytical Boltzmann weights.
Table S1: Summary Müller-Brown Summary of the iterative procedure for Müller-Brown potential.
@@
Table S2: Summary Alanine Summary of the iterative procedure for Alanine.
@@
Figure S3: Iterations sampling Alanine Sampling at different iterations on the procedure on Alanine. Top row: Scatter plot of the sampled points in the $\varphi\psi$ space. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red.. Middle row: Time series of the $\varphi$ value from a trajectory started from state A. Bottom row: Time series of the $\varphi$ value from a trajectory started from state B.
@@
Table S3: Summary double-path potential Summary of the iterative procedure for double path potential.
@@
Figure S4: Iterations sampling double-path Sampling at different iterations of the procedure on the double path potential. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red. The 0.5 value of the committor estimate at that iteration is provided as a black line.
@@
Table S4: Summary Chignolin Summary of the iterative procedure for chignolin.
@@
Comparison between different descriptor sets In Ref. 3, we trained the committor model using as inputs the 45 distances between $\alpha$-C atoms, whereas here, we used the same set of 210 distances identified in Ref. 19. To compare the two descriptor sets, we optimized the committor model on the same dataset using the two sets, obtaining a $K_{m}$ value of 2.17 for the 210 distances and 18.5 for the 45 distances. This result clearly shows how the greater variational flexibility introduced by side-chain information allows for capturing more features of the folding process.
@@
Table S5: LASSO summary Chignolin Summary of LASSO analysis for chignolin. The atoms involved in the distances used as descriptors are indicated according to the indexes used in PLUMED (upper table) and their names from the PDB (lower table).
@@
Figure S5: Free energy surface Chignolin Free energy estimates for Chignolin projected along $z$ and $q$ obtained averaging on 6 independent simulations: 3 started from the folded state and 3 from the unfolded one. The average value and the relative standard deviation are reported as a solid line and a shaded region, respectively.
@@
The funnel restraint In our simulations, we used a funnel restraint Limongelli, Bonomi, and Parrinello (2013) equivalent to the one previously employed by Refs.38; 70; 71 on the same system. Here, we summarize the details of such a restraint, while more details can be found in the original works. The funnel limits the space available to the ligand in the unbound state U by confining it to a cylindrical volume above the binding site. As the ligand approaches the binding site, the funnel restraint becomes wider so that its presence does not affect the binding process itself. Having aligned with PLUMED the system to a reference configuration where the binding axis is found along the vertical axis, we define $h$ as the projection on the binding axis of the center of the carbon atoms of each ligand and $r$ as its radial component. When $h>10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel surface is a cylinder with radius $R_{cyl}=2$\mathrm{\SIUnitSymbolAngstrom}$$ with its axis along the vertical direction. When $h<10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel opens into an umbrella-like shape with a 45 degree angle whose surface is defined by $r=12-h$. The force that, for a displacement x, pushes the ligand away from the funnel’s surface is harmonic $-k_{F}x$ with $kF=20$ kJ/mol Å$^{-2}$. A further harmonic restraint is applied on $h$ to prevent the ligand from getting too far from the host, reaching the upper boundary of the simulation box. The corresponding force is $-k_{U}(h-18)$ for $h>18$\mathrm{\SIUnitSymbolAngstrom}$$ and $k_{U}=40$ kJ/mol Å$^{-2}$.
@@
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref. 38, 8 with respect to 2.5Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)), and a set of distances between the same 4 ligand atoms and the binding pocket, instead of the $h$ projection only, as inputs of a neural network (NN) with architecture [16, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for $\sim$30000 epochs. The $\alpha$ hyperparameter in the loss function was set to 1. In the loss function, we assigned to the virtual points along the binding axis used for the computation of water coordination numbers an infinite mass, that is, $10^{6}$. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S6 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S6: Summary calixarene Summary of the iterative procedure for calixarene.
@@
Figure S6: Feature relevance calixarene Descriptors ranking for the committor model of calixarene trained using the descriptors listed in the computational details.
@@
Figure S7: 1D free energy surface calixarene Free energy estimates for calixarene projected along the projection of the ligand molecule on the binding axis $h$, $z$ and $q$.
Figure S8: 2D free energy surface calixarene Two-dimensional free energy surface for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
@@
Figure S9: Kolomogorov distribution calixarene Two-dimensional Kolmogorov probability $p_{K}$ projection for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).